1. Our company reported the following financial numbers for one of its divisions for the year; average total assets of $4,100,000; sales of $4,525,000; cost of goods sold of $2,550,000; and operating expenses of $1,372,000. Assume a target income of 10% of average invested assets. The investment turnover is:

2. A company's flexible budget for 12,000 units of production showed sales, $48,000; variable costs, $18,000; and fixed costs, $16,000. The sales expected if the company produces and sells 16,000 units is:
Answer #1

1. Investment turnover = Sales / average total assets

= 4525000/4100000

= 1.10

2. Sales at 16000 units = 48000*16000/12000 = $64000